# Probing supermassive black hole scalarization with Pulsar Timing Arrays

### v1

@@ III.2 Characteristic strain in GR and EsGB gravity
where we omitted the dependence on the redshift and source parameters for simplicity. The quantity ${\rm d}E_{\text{gw}}^{s}/{\rm d}t_{s}$ is the instantaneous GW power emitted in GWs in the source rest-frame. From App. A, it is given in EsGB gravity at lowest post-Newtonian (PN) order by
@@
As specified below Eq. (14), the underlying astrophysical model plays a role in setting the normalization of $h_{c}$, while the theory of gravity (GR in this case) defines the frequency scaling.
Let us now consider EsGB gravity. In this case, the frequency evolution is modified due to an additional energy loss of the binary in the scalar sector, so that (see App. A):
@@ III.3 Qualitative analysis of the spectrum
where $\chi=\{\log_{10}\mathcal{M},\log_{10}q\}$. As expected, the evolution in the early inspiral (dotted blue line) is driven by the dipolar scalar emission, which allows for a faster orbital shrinking than GR. At the turnover frequency, defined by Eq. (26) and plotted as a red dashed vertical line, the quadrupolar GW emission becomes equal to the dipolar scalar emission, and eventually takes over at larger frequencies. The resulting characteristic strain (green line), produced using Eq. (III.2), smoothly interpolates between these two different regimes. As a consequence, the spectral index varies continuously from $1/3$ at low frequencies to $2/3$, at the high end of the PTA band. We notice that such a behavior is compatible with recent observations. Indeed, for the parameters used in Fig. 2, EsGB gravity predicts a flatter spectral index than GR at low frequencies. This matches what has been recently observed by PTA collaborations: as an example, in Fig. 3, produced using the EPTA DR2New dataset [70], we see that the scaling predicted by EsGB in the dipolar-emission regime (black dotted line), as computed by Eq. (23), is visually more in agreement with the data compared to the usual scaling predicted by circularly inspiralling SMBHBs in GR (red dashed line). To produce this plot, we expressed the power spectrum of timing residuals in Eq. (13) as a power-law, namely
@@
However, we must stress that there is a number of different astrophysical effects that could explain a flattening of the spectral index, such as high eccentricities or interaction with the environment [31, 32, 33, 34, 35, 36, 37]. Moreover, when computing the characteristic strain produced by a realistic population of SMBHBs, the nice picture developed above becomes more fuzzy, as different binaries generally have different turnover frequencies. In any case, the flattening of the spectral index still remains a solid prediction of the EsGB theory, which, for the SGWB under examination, generally yields a value in the range $n\in[1/3,2/3]$. Overall, no conclusive statement can be made at the moment. Future data are expected to improve our understanding of environmental and astrophysical effects and may lead to a final answer.
@@ IV.1 Mock injection
We consider a future PTA experiment, with a timespan of $T_{\text{obs}}=30\,\text{yr}$. This sets the minimum frequency and the width of the frequency bins to $f_{l}=\Delta f\sim 1/T_{\text{obs}}$. Moreover, we assume we are able to measure the SGWB in the first 20 frequency bins. While this is an idealized setup, it serves to provide a clear illustration of our analysis in a simplified framework. Then, we compute the characteristic strain produced by the astrophysical population in Eq. (29), fixing the overall normalization to $\log_{10}(A/\text{Mpc}^{-3})=-4$ and injecting a non-zero value of $\log_{10}\lambda_{9}$ in Eq. (14), namely $\log_{10}\lambda_{9}=1.9$. In constructing $h_{c}(f)$, we integrate the chirp mass in the mass range $\log_{10}\mathcal{M}/\text{M}_{\odot}\in[7,10.5]$. Chirp masses smaller than the low mass end are expected to contribute negligibly to the signal [63], while the high mass end is chosen so as not to produce too massive BHs (for $\mathcal{M}\sim 10^{10.5}\,\text{M}_{\odot}$ and $\log_{10}q\sim-1.5$, the mass of the primary is $m_{1}\sim 2.5\cdot 10^{11}\,\text{M}_{\odot}$).
@@
[…]
@@
In the last line, we denoted the uncertainty on the measured value of $\mu_{i}$ by $\sigma_{i}=\epsilon\mu_{i}$. The case $\epsilon=1/20$ may realistically be achieved in the future for some frequency bins; the other two scenarios are highly idealized (see e.g. Ref. [27, 22]). Nonetheless, as mentioned above, they serve to illustrate our analysis in a simplified setting. As priors, we use uninformative ranges, namely $\log_{10}(A/\text{Mpc}^{-3})\in[-7,0]$ and $\log_{10}\lambda_{9}=[-5,4]$, so as not to bias the recovery.
In Fig. 4, we plot the result of the Bayesian inference for $\epsilon=1/100$, while we display the results for $\epsilon=1/50,1/20$ in App. B. We find that the amplitude of the injected merger density is quite well constrained. In contrast, while the parameter $\log_{10}\lambda_{9}$ is clearly detected in the data, its posterior distribution is bimodal and spans several orders of magnitude—even within this idealized scenario. However, the shape of the 2D corner plot elucidates the nature of the degeneracy and, to some extent, clarifies why we are fundamentally limited in precisely extracting the true value of $\lambda_{9}$.
@@
Indeed, let us remark that, in the theory we are considering, a given value of $\lambda_{9}$ predicts a window of scalarized masses, as shown in Fig. 1. As $\lambda_{9}$ increases (decreases), the window shifts to higher (lower) BH masses, maintaining its amplitude and width (in log space) constant. Therefore, considering a system of two SMBHBs with different masses, there is a degeneracy: in other words, as the frequency evolution in the dipolar-dominated regime is only concerned with the absolute difference of scalar charges (at least at lowest PN order, see Eq. (III.2)), we cannot distinguish the situation in which the window is centered on one of the two masses from the opposite one. As a consequence, the constraints on $\log_{10}\lambda_{9}$ will manifest this degeneracy, as clearly shown in Fig. 4. As a matter of fact, such a degeneracy can be partially broken by the astrophysical population. To gain more insight, let us start by rewriting Eq. (14) as
@@
Figure 5: Scalarized window of SMBH masses and source loudness, as defined in Eq. (33) for $f=10^{-9}\text{Hz}$. The curves are produced assuming a uniform distribution in $\log_{10}q$ ranging from $\log_{10}q\in[-1.5,0]$, and assuming the merger density in Eq. (29). The top panel displays the scalarization window for $\log_{10}\lambda_{9}=1.9$ and the resulting source loudness in EsGB theory. The middle and low panels show the same curves, but for $\log_{10}\lambda_{9}=1.6$ and $\log_{10}\lambda_{9}=0.9$ In the three panels, we also overplot the source loudness in GR, which does not depend neither on the scalarization window nor on the mass ratios of the systems.
@@
Finally, we observe the presence of a valley between the two peaks of the bimodal distribution in Fig. 4. As shown in the middle panel of Fig. 5, this feature arises because, in that region of parameter space, most SMBHB systems consist of two scalarized black holes with similar values of the adimensional charge $\alpha$. In such cases, the charge difference is generally small, leading to weaker dipolar emission compared to scenarios where only one of the two black holes is scalarized. Consequently, the peak of the source loudness in EsGB gravity is not significantly modified compared to GR, implying that the overall shape of the SGWB remains largely unchanged. This results in a suppression of the posterior in the corresponding region of $\log_{10}\lambda_{9}$. Nonetheless, a non-negligible number of posterior samples still populate this region, and they tend to be correlated with a lower amplitude for the merger rate. This is ultimately caused by the fact that, as we have shown in Figs. 2 and 5, the presence of EsGB gravity suppresses the overall amplitude of the SGWB. Therefore, the mock data points we constructed can also be fitted - although with a smaller likelihood - by a smaller $\log_{10}(A/\text{Mpc}^{-3})$ and values of $\log_{10}\lambda_{9}$ which do not produce a heavy suppression of the SGWB. As a consequence, our analysis cannot fully exclude the values of $\log_{10}\lambda_{9}$ in the trough of Fig. 4, although strongly disfavoring them.
The last point becomes increasingly relevant as the precision of our mock measurements decreases - that is, if $\epsilon$ increases. In App. B (Fig. 7), we show the results for $\epsilon=1/50,1/20$. In the former case, the GR fit is more in agreement with the data compared to the $\epsilon=1/100$ case: the posterior in $\log_{10}(A/\text{Mpc}^{-3})$ becomes bimodal - signaling that the data could be fitted by pure GR, although with a smaller likelihood - while the through in $\log_{10}\lambda_{9}$ disappears. In the latter case, the uncertainty in the mock measurements is so large that even the trough in the amplitude disappears, resulting in significantly broader constraints. However, the characteristic structure in the 2D corner plot remains discernible and continues to serve as a distinctive imprint of EsGB gravity on the data.
@@ IV.2 Application to real data
where $\text{Prob}_{i}$ is the normalized probability distribution constructed from the $i$-th bin in Fig. 1 of Ref. [27]. We consider $N_{\text{bins}}=9$, as commonly done in the literature. However, we point out that this might introduce some bias in the analysis, as spikes from single sources dominating over the background are expected to show up at frequencies larger than $f\gtrsim 10^{-8}\,\text{Hz}$ [63]. In this case, the parametric form in Eq. (14) with the merger rate distribution in Eq. (15) is not expected to hold anymore, and the distribution of the resulting characteristic strain around the value in Eq. (14) may exhibit significant scatter, due to the particular realization of the Universe [71]. To account for this, in App. B (Fig. 8) we present an analysis performed on the first 4 frequency bins, which are expected to be well within the region of validity of the description in Eqs. (14) and (15). As done in Sec. IV.1, we integrate the chirp mass in the range $\log_{10}\mathcal{M}/\text{M}_{\odot}\in[7,10.5]$ and the redshift $z$ in the range $z\in[0,5]$. The priors on the parameters of the search are taken from Sec. III.1, and are chosen to be consistent with Refs. [66, 27]. Furthermore, we sample $\log_{10}\lambda_{9}$ in the range $\log_{10}\lambda_{9}\in[-5,4]$, which largely encompasses all the possible values of $\lambda_{9}$ that may scalarize SMBHBs in the PTA mass range. We present the results of the analysis in Fig. 6, while we show the results of a similar analysis of the first 14 frequency bins the NANOGrav 15yr dataset [72, 22] to Fig. 9 in App. B. The amplitude of the comoving merger density ($\log_{10}\frac{\dot{n}_{0}}{\text{Mpc}^{-3}\text{Gyr}^{-1}}$) is well constrained, whereas the remaining parameters largely reflect the prior. We report a mild excess in the posterior of $\log_{10}\lambda_{9}$ in the region $-2\lesssim\log_{10}\lambda_{9}\lesssim 0$, which implies a slight preference for $1.5\cdot 10^{7}~{}\text{km}\lesssim\lambda\lesssim 1.5\cdot 10^{9}~{}\text{km}$. However, current PTA data are not sensitive enough yet to conclusively discriminate between the presence of EsGB gravity and GR. This is reflected in the marginalized posterior of $\log_{10}\lambda_{9}$, which more closely resembles the right panel of Fig. 7 in the Appendix than Fig. 4. Qualitatively, values of $\log_{10}\lambda_{9}$ sufficiently above or below the peak yield a SGWB spectrum that is indistinguishable from that predicted by GR, as the scalarization window shifts to masses outside the range contributing to the PTA signal (see also Fig. 5 for a visual comparison). This ultimately explains why, if the data are not precise enough to disentangle EsGB gravity from GR, the posterior for $\log_{10}\lambda_{9}$ shows no fall-off behavior at either large or small values (as opposed to what happens, for instance, in the idealized case of Fig. 4). The enhancement is even less pronounced in the 4-bins analysis in Fig. 8 in App. B, and, in general, the posteriors remain too broad to allow for the identification of a distinctive shape in the 2D corner plot of $\{\log_{10}\frac{\dot{n}_{0}}{\text{Mpc}^{-3}\text{Gyr}^{-1}},\log_{10}\lambda_{9}\}$, such as the one in Fig. 4. Improved data quality and a better understanding of the underlying astrophysical SMBHB population could help determine whether this excess is a mere statistical fluctuation or a genuine hint for a fundamental physical process. In Fig. 6, we examine two reference distributions for the mass ratio $q$: a log-uniform distribution over $\log_{10}q\in[-1.5,0]$ (solid black lines) and a uniform distribution in the same range, i.e. $q\in[0.03,1]$ (solid red lines). The resulting posteriors do not show any appreciable difference, and we present only the former choice in Figs. 8 and 9 in App. B. This qualitatively illustrates that our findings are fairly robust against variations in the mass ratio distribution, provided they are not strongly peaked around $q=1$, in which case the effect of EsGB gravity would be suppressed, as already pointed out in Sec. III.1.
@@
Should the peak in the $\log_{10}\lambda_{9}$ posterior be confirmed, it would still be essential to rule out alternative explanations, such as strong eccentricity or significant dynamical friction, before interpreting it as evidence for EsGB gravity. As discussed earlier, these effects are largely degenerate. Future measurements may help break this degeneracy and clarify the origin of the observed features.

### v2

@@ III.2 Characteristic strain in GR and EsGB gravity
where we omitted the dependence on the redshift and source parameters for simplicity. The quantity ${\rm d}E_{\text{gw}}^{s}/{\rm d}t_{s}$ is the instantaneous GW power emitted in GWs in the source rest-frame. From Appendix A, it is given in EsGB gravity at lowest post-Newtonian (PN) order by
@@
As specified below Eq. (14), the underlying astrophysical model plays a role in setting the normalization of $h_{c}$, while the theory of gravity (GR in this case) defines the frequency scaling, characterized by a spectral index $n=2/3$.
Let us now consider EsGB gravity. In this case, the frequency evolution is modified due to an additional energy loss of the binary in the scalar sector, so that (see Appendix A):
@@ III.3 Qualitative analysis of the spectrum
where $\chi=\{\log_{10}\mathcal{M},\log_{10}q\}$. As expected, the evolution in the early inspiral (dotted blue line) is driven by the dipolar scalar emission, which allows for a faster orbital shrinking than GR. At the turnover frequency, defined by Eq. (26) and plotted as a red dashed vertical line, the quadrupolar GW emission becomes equal to the dipolar scalar emission, and eventually takes over at larger frequencies. The resulting characteristic strain (green line), produced using Eq. (III.2), smoothly interpolates between these two different regimes. As a consequence, the spectral index varies continuously from $1/3$ at low frequencies to $2/3$, at the high end of the PTA band.
We notice that such a behavior is compatible with recent observations. Indeed, for the parameters used in Fig. 2, EsGB gravity predicts a flatter spectral index than GR at low frequencies. This matches what has been recently observed by PTA collaborations: as an example, in Fig. 3, produced using the EPTA DR2New dataset [71], we see that the scaling predicted by EsGB in the dipolar-emission regime (black dotted line), as computed by Eq. (23), is visually more in agreement with the data compared to the usual scaling predicted by circularly inspiralling SMBHBs in GR (red dashed line). To produce this plot, we expressed the power spectrum of timing residuals in Eq. (13) as a power-law, namely
@@
However, we must stress that there is a number of different astrophysical effects that could explain a flattening of the spectral index, such as high eccentricities or interaction with the environment [31, 32, 33, 34, 35, 36, 37]. While a complete and quantitative comparison among these effects lies beyond the purpose of this work, and could well be the subject of a dedicated analysis, we point out that, in principle, they can still be distinguishable. To illustrate this, we follow the analysis of Ref. [34], which examines orbital shrinking driven by stellar interactions or binary eccentricity. For a circular binary with chirp mass $\mathcal{M}$ and mass ratio $q$, whose inspiral is driven by interactions with stars, the residence time in a log-frequency bin reads
$\frac{{\rm d}t_{s}}{{\rm d}\log f_{s}}=\frac{2}{3}\frac{\sigma}{H\rho_{i}}\frac{q^{1/5}}{(1+q)^{2/5}}\frac{x}{G^{2}\mathcal{M}},$ (29)
where $H\sim 15\div 20$ is a dimensionless constant, $\sigma\sim 200~\text{km}/\text{s}$ is the typical stellar-bulge velocity dispersion and $\rho_{i}\sim 100~\text{M}_{\odot}\text{pc}^{-3}$ is the mean stellar density at the binary influence radius, i.e. the radius encompassing a stellar mass twice the total binary mass [34] [^5] Notice the difference in the prefactor compared to Eq. 19 in Ref. [34]. This arises because, while Ref. [34] considers the orbital frequency evolution, here we focus on the GW emission frequency, which for circular binaries is twice the orbital frequency. Combining it with Eq. (18) (now setting $\alpha_{A},\alpha_{B}=0$) and substituting the result in Eq. (14) gives
$h_{c}^{2}(f)=A_{\text{star}}f^{2},$ (30)
where we omit the explicit form of $A_{\text{star}}$ for simplicity. Therefore, if the SGWB is dominated by binaries whose orbital evolution is driven by interactions with stars, the spectral index is predicted to be 1-1. However, at sufficiently high frequencies, GW emission is expected to be the primary source of orbital shrinking. Consequently, the shape of the characteristic strain will feature a peak at a frequency $f_{\text{peak}}$, which can be determined by comparing Eq. (29) with Eq. (19). Depending on how strong the interaction with the environment is, then, the spectral index can assume any value between [1,2/3][-1,2/3].
On the other hand, as we have shown in Sec. III.3 and in Fig. 2, the EsGB theory of gravity can only yield a spectral index within the range $[1/3,2/3]$, with the lower bound corresponding to the case where all binaries contributing to the SGWB evolve primarily through EsGB dipolar emission (see Eq. (23)). Therefore, if observational data pointed at the presence of a peak in the characteristic strain, rather than a modest flattening of the spectral index, the EsGB hypothesis would be strongly disfavored compared to models involving stellar interactions.
Binary eccentricity is also expected to generate a peak in the characteristic strain. However, as eccentric binaries emit across different frequency bins, calculating the characteristic strain at a specific observing frequency $f$ is less straightforward. Nonetheless, for the purposes of this analysis, the main conclusions remain consistent with those drawn for stellar interactions. We refer the reader to Ref. [34] for a comprehensive treatment of binary eccentricity.
Thus, from a theoretical standpoint, it is possible to distinguish between astrophysical effects - such as interactions with stars and binary eccentricity - and modifications of gravity within the EsGB framework. However, the problem becomes more intricate when computing the characteristic strain produced by a realistic population of SMBHBs: in the EsGB theory, different binaries generally have different turnover frequencies; similarly, for interactions with stars or eccentricity, the location of the peak frequency depends on the underlying properties of the astrophysical population of SMBHBs and their environment. In any case, the flattening of the spectral index still remains a solid prediction of the EsGB theory, which, for the SGWB under examination, generally yields a value in the range $n\in[1/3,2/3]$. Overall, no conclusive statement can be made at the moment. Future data are expected to improve our understanding of environmental and astrophysical effects and may lead to a final answer.
@@ IV.1 Mock injection
We consider a future PTA experiment, with a timespan of $T_{\text{obs}}=30\,\text{yr}$. This sets the minimum frequency and the width of the frequency bins to $f_{l}=\Delta f\sim 1/T_{\text{obs}}$. Moreover, we assume we are able to measure the SGWB in the first 20 frequency bins. While this is an idealized setup, it serves to provide a clear illustration of our analysis in a simplified framework. Then, we compute the characteristic strain produced by the astrophysical population in Eq. (31), fixing the overall normalization to $\log_{10}(A/\text{Mpc}^{-3})=-4$ and injecting a non-zero value of $\log_{10}\lambda_{9}$ in Eq. (14), namely $\log_{10}\lambda_{9}=1.9$. In constructing $h_{c}(f)$, we integrate the chirp mass in the mass range $\log_{10}\mathcal{M}/\text{M}_{\odot}\in[7,10.5]$. Chirp masses smaller than the low mass end are expected to contribute negligibly to the signal [64], while the high mass end is chosen so as not to produce too massive BHs (for $\mathcal{M}\sim 10^{10.5}\,\text{M}_{\odot}$ and $\log_{10}q\sim-1.5$, the mass of the primary is $m_{1}\sim 2.5\times 10^{11}\,\text{M}_{\odot}$).
@@
$\begin{split}\delta t_{\text{rms},i}&\sim\left(\int_{\Delta f}P(f)df\right)^{1/2}\sim(P(f_{i})\Delta f)^{1/2}\\ &\sim\left(\frac{P(f_{i})}{T_{\text{obs}}}\right)^{1/2}.\end{split}$ (32)
@@
In the last line, we denoted the uncertainty on the measured value of $\mu_{i}$ by $\sigma_{i}=\epsilon\mu_{i}$. The case $\epsilon=1/20$ may realistically be achieved in the future for some frequency bins; the other two scenarios are highly idealized (see e.g. Refs. [27, 22]). Nonetheless, as mentioned above, they serve to illustrate our analysis in a simplified setting. As priors, we use uninformative ranges, namely $\log_{10}(A/\text{Mpc}^{-3})\in[-7,0]$ and $\log_{10}\lambda_{9}=[-5,4]$, so as not to bias the recovery.
In Fig. 4, we plot the result of the Bayesian inference for $\epsilon=1/100$, while we display the results for $\epsilon=1/50,1/20$ in Appendix C. We find that the amplitude of the injected merger density is quite well constrained. In contrast, while the parameter $\log_{10}\lambda_{9}$ is clearly detected in the data, its posterior distribution is bimodal and spans several orders of magnitude—even within this idealized scenario.
However, the shape of the 2D corner plot elucidates the nature of the degeneracy and, to some extent, clarifies why we are fundamentally limited in precisely extracting the true value of $\lambda_{9}$.
@@
Indeed, let us remark that, in the theory we are considering, a given value of $\lambda_{9}$ predicts a window of scalarized masses, as shown in Fig. 1. As $\lambda_{9}$ increases (decreases), the window shifts to higher (lower) BH masses, maintaining its amplitude and width (in log space) constant. Therefore, considering a system of two SMBHBs with different masses, there is a degeneracy: in other words, as the frequency evolution in the dipolar-dominated regime is only concerned with the absolute difference of scalar charges (at least at lowest PN order, see Eq. (III.2)), we cannot distinguish the situation in which the window is centered on one of the two masses from the opposite one. As a consequence, the constraints on $\log_{10}\lambda_{9}$ will manifest this degeneracy, as clearly shown in Fig. 4.
As a matter of fact, such a degeneracy can be partially broken by the astrophysical population. To gain more insight, let us start by rewriting Eq. (14) as
@@
Figure 5: Scalarized window of SMBH masses and source loudness, as defined in Eq. (35) for $f=10^{-9}\text{Hz}$. The curves are produced assuming a uniform distribution in $\log_{10}q$ ranging from $\log_{10}q\in[-1.5,0]$, and assuming the merger density in Eq. (31). The top panel displays the scalarization window for $\log_{10}\lambda_{9}=1.9$ and the resulting source loudness in EsGB theory. The middle and low panels show the same curves, but for $\log_{10}\lambda_{9}=1.6$ and $\log_{10}\lambda_{9}=0.9$ . In the three panels, we also overplot the source loudness in GR, which does not depend neither on the scalarization window nor on the mass ratios of the systems.
@@
Finally, we observe the presence of a valley between the two peaks of the bimodal distribution in Fig. 4. As shown in the middle panel of Fig. 5, this feature arises because, in that region of parameter space, most SMBHB systems consist of two scalarized black holes with similar values of the adimensional charge $\alpha$. In such cases, the charge difference is generally small, leading to weaker dipolar emission compared to scenarios where only one of the two black holes is scalarized. Consequently, the peak of the source loudness in EsGB gravity is not significantly modified compared to GR, implying that the overall shape of the SGWB remains largely unchanged. This results in a suppression of the posterior in the corresponding region of $\log_{10}\lambda_{9}$.
Nonetheless, a non-negligible number of posterior samples still populate this region, and they tend to be correlated with a lower amplitude for the merger rate. This is ultimately caused by the fact that, as we have shown in Figs. 2 and 5, the presence of EsGB gravity suppresses the overall amplitude of the SGWB. Therefore, the mock data points we constructed can also be fitted - although with a smaller likelihood - by a smaller $\log_{10}(A/\text{Mpc}^{-3})$ and values of $\log_{10}\lambda_{9}$ which do not produce a heavy suppression of the SGWB. As a consequence, our analysis cannot fully exclude the values of $\log_{10}\lambda_{9}$ in the trough of Fig. 4, although strongly disfavoring them.
The last point becomes increasingly relevant as the precision of our mock measurements decreases - that is, if $\epsilon$ increases. In Appendix C (Fig. 8), we show the results for $\epsilon=1/50,1/20$. In the former case, the GR fit is more in agreement with the data compared to the $\epsilon=1/100$ case: the posterior in $\log_{10}(A/\text{Mpc}^{-3})$ becomes bimodal - signaling that the data could be fitted by pure GR, although with a smaller likelihood - while the through in $\log_{10}\lambda_{9}$ disappears. In the latter case, the uncertainty in the mock measurements is so large that even the trough in the amplitude disappears, resulting in significantly broader constraints. However, the characteristic structure in the 2D corner plot remains discernible and continues to serve as a distinctive imprint of EsGB gravity on the data.
@@ IV.2 Application to real data
where $\text{Prob}_{i}$ is the normalized probability distribution constructed from the $i$-th bin in Fig. 1 of Ref. [27]. We consider $N_{\text{bins}}=9$, as commonly done in the literature. However, we point out that this might introduce some bias in the analysis, as spikes from single sources dominating over the background are expected to show up at frequencies larger than $f\gtrsim 10^{-8}\,\text{Hz}$ [64]. In this case, the parametric form in Eq. (14) with the merger rate distribution in Eq. (15) is not expected to hold anymore, and the distribution of the resulting characteristic strain around the value in Eq. (14) may exhibit significant scatter, due to the particular realization of the Universe [72]. To account for this, in Appendix C (Fig. 9) we present an analysis performed on the first four frequency bins, which are expected to be well within the region of validity of the description in Eqs. (14) and (15).
As done in Sec. IV.1, we integrate the chirp mass in the range $\log_{10}\mathcal{M}/\text{M}_{\odot}\in[7,10.5]$ and the redshift $z$ in the range $z\in[0,5]$. The priors on the parameters of the search are taken from Sec. III.1, and are chosen to be consistent with Refs. [67, 27]. Furthermore, we sample $\log_{10}\lambda_{9}$ in the range $\log_{10}\lambda_{9}\in[-5,4]$, which largely encompasses all the possible values of $\lambda_{9}$ that may scalarize SMBHBs in the PTA mass range. We present the results of the analysis in Fig. 6, while we show the results of a similar analysis of the first 14 frequency bins the NANOGrav 15yr dataset [73, 22, 74] to Fig. 10 in Appendix C.
The amplitude of the comoving merger density ($\log_{10}\frac{\dot{n}_{0}}{\text{Mpc}^{-3}\text{Gyr}^{-1}}$) is well constrained, whereas the remaining parameters largely reflect the prior. We report a mild excess in the posterior of $\log_{10}\lambda_{9}$ in the region $2\lesssim\log_{10}\lambda_{9}\lesssim 0$, which implies a slight preference for $1.5\times 10^{7}~\text{km}\lesssim\lambda\lesssim 1.5\times 10^{9}~\text{km}$. However, the Bayes factor comparing EsGB to GR is only $\sim 1.35$, suggesting that this preference is not statistically significant (see Appendix B for details on the Bayes factor calculation). Therefore, although we report a small indication for $1.5\times 10^{7}~\text{km}\lesssim\lambda\lesssim 1.5\times 10^{9}~\text{km}$, current PTA data are not sensitive enough yet to conclusively discriminate between the presence of EsGB gravity and GR. This is reflected in the marginalized posterior of $\log_{10}\lambda_{9}$, which more closely resembles the right panel of Fig. 8 in the Appendix than Fig. 4.
Qualitatively, values of $\log_{10}\lambda_{9}$ sufficiently above or below the peak yield a SGWB spectrum that is indistinguishable from that predicted by GR, as the scalarization window shifts to masses outside the range contributing to the PTA signal (see also Fig. 5 for a visual comparison). This ultimately explains why, if the data are not precise enough to disentangle EsGB gravity from GR, the posterior for $\log_{10}\lambda_{9}$ shows no fall-off behavior at either large or small values (as opposed to what happens, for instance, in the idealized case of Fig. 4).
The enhancement is even less pronounced in the 4-bins analysis in Fig. 9 in Appendix C (Bayes factor $\sim 1.25$), and, in general, the posteriors remain too broad to allow for the identification of a distinctive shape in the 2D corner plot of $\{\log_{10}\frac{\dot{n}_{0}}{\text{Mpc}^{-3}\text{Gyr}^{-1}},\log_{10}\lambda_{9}\}$, such as the one in Fig. 4. Improved data quality and a better understanding of the underlying astrophysical SMBHB population could help determine whether this excess is a mere statistical fluctuation or a genuine hint for a fundamental physical process.
In Fig. 6, we examine two reference distributions for the mass ratio $q$: a log-uniform distribution over $\log_{10}q\in[-1.5,0]$ (solid black lines) and a uniform distribution in the same range, i.e. $q\in[0.03,1]$ (solid red lines). The resulting posteriors do not show any appreciable difference, and we present only the former choice in Figs. 9 and 10 in Appendix C. This qualitatively illustrates that our findings are fairly robust against variations in the mass ratio distribution, provided they are not strongly peaked around $q=1$, in which case the effect of EsGB gravity would be suppressed, as already pointed out in Sec. III.1.
@@
Should the peak in the $\log_{10}\lambda_{9}$ posterior be confirmed, it would still be essential to rule out alternative explanations, such as strong eccentricity or significant interaction with the environment, before interpreting it as evidence for EsGB gravity. As discussed earlier, while these effects present characteristic distinguishable signatures, they may be degenerate to some extent when considering a realistic SMBHB population. Future measurements may help break this degeneracy and clarify the origin of the observed features.